x^5-3x^3-4x=0
solve under the set of complex numbers

Answer by Alan3354(69443)

x^5-3x^3-4x=0
solve under the set of complex numbers
------------------
x = 0 is one.
----
x^4 - 3x^2 - 4 = 0
(x^2 - 4)*(x^2 + 1) = 0
x = +2, -2
x = +i, -i